OpenStudy (anonymous):

For f(x)=4x+1 and g(x)=x^2-5 , find (f-g) (x) . i think the answer is -x^2+4x-4

OpenStudy (whpalmer4):

check your algebra again...

OpenStudy (whpalmer4):

4x+1 - (x^2 - 5) = 4x+1 - x^2 - (-5) =

OpenStudy (whpalmer4):

have to be careful when subtracting polynomials like this — I prefer to multiply the second polynomial by -1 first and add...seem to make fewer mistakes that way. your mileage may vary

OpenStudy (anonymous):

so x^2-4x-6?

OpenStudy (whpalmer4):

no... 4x + 1 - (x^2 - 5) = 4x + 1 - x^2 - (-5) subtracting a negative number is the same as adding a positive number of the same magnitude

OpenStudy (anonymous):

no +6*

OpenStudy (whpalmer4):

you're being way too sloppy here with the signs :-( 4x + 1 - (x^2 - 5) = 4x + 1 - x^2 - (-5) = 4x + 1 - x^2 + 5 = -x^2 + 4x + 6

OpenStudy (anonymous):

i see.. thank you :)